OpenStudy (anonymous):

A website that sells songs for downloading increased its price per song from $0.99 to $1.29. Ariana spent $15.36 downloading songs before and after the price increase. She downloaded 4 more songs at $0.99 than at $1.29. The set of equations below represents the situation where x is the number of songs Ariana downloaded at $0.99 and y is the number of songs she downloaded at $1.29. x = y + 4 .99x + 1.29y = 15.36 What is the exact number of songs Ariana downloaded at the $0.99 price?

OpenStudy (anonymous):

@johnweldon1993 would you mind helping me again?

OpenStudy (johnweldon1993):

Well for this, we have a system of equations and we just want to solve for 'x'

OpenStudy (johnweldon1993):

Look at that top equation \[\large x = y + 4\] We want to solve this for 'y'...so that way, when we substitute it into the second equation, we will only have 'x' to deal with so solve that for 'y'

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

\[x = y + 4\]\[-4x = y \]

OpenStudy (anonymous):

@johnweldon1993 are you still on